# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ 4.2 Proofs of the Error Bound
which implies $\|\bm{\Sigma}_{L}^{(i)}-\sigma_{t_{i}}^{*}\bm{I}\|\leq\sigma_{\min}^{*}/2$. This, together with Lemma 5 and $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$, yields
$\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}\geq\sigma_{\min}^{*}\|\bm{T}_{l}^{(i)}\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{L}^{(i)}\bm{T}_{l}^{(i)}\|_{F}.$
Using this and (138), we obtain (134). ∎
Since the assumptions are different for the cases $L=2$ and $L\geq 3$ to bound the singular values by the gradient norm, we prove 5 by addressing these two scenarios separately. Each case requires a distinct approach to account for the structural differences due to the varying depth of the networks.
Proof of 5.
@@
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. According to (77a) or (79), we have $\mathrm{dist}(\bm{W},\mathcal{W})\leq\delta_{\sigma}/3<\sigma_{\min}^{*}/2$. This, together with Lemma 3 and 2, implies that (57)-(60) and (63) hold. According to $\mathrm{dist}(\bm{W},\mathcal{W})\leq\delta_{\sigma}/3$, 4, and Lemma 4, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for all $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (68), (69), and (70) hold for all $l=2,\dots,L$, where $\bm{A}_{i}$ and $\bm{B}_{i}$ for each $i\in[p+1]$ are respectively defined in (71) and (72). Recall that (66) denotes an SVD of $\bm{W}_{l}$ and let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ and $\bm{T}_{l}:=\mathrm{BlkDiag}(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)})$ for each $l=2,\dots,L$. This, together with (68), implies
@@
(i) For ease of exposition, let $\bm{\Psi}:=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}\in\mathbb{R}^{d_{2}\times d_{0}}$. Now, we partition $\bm{\Psi}$ into the following block form
@@
where $\bm{\Psi}^{(i,j)}\in\mathbb{R}^{g_{i}\times g_{j}}$ for each $i,j\in[p]$, and $\bm{\Psi}^{(p+1,j)}\in\mathbb{R}^{(d_{2}-r_{\sigma})\times g_{j}},\ \bm{\Psi}^{(i,p+1)}\in\mathbb{R}^{g_{i}\times(d_{0}-r_{\sigma})}$ for each $i,j\in[p]$, and $\bm{\Psi}^{(p+1,p+1)}\in\mathbb{R}^{(d_{2}-r_{\sigma})\times(d_{0}-r_{\sigma})}$. According to (69), we have
$c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\left\|(\bm{\Sigma}_{1}\bm{\Sigma}_{1}^{T})\bm{\Sigma}_{1}+\lambda\bm{\Sigma}_{1}-\sqrt{\lambda}\mathrm{BlkDiag}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{A}_{p+1}\right)\bm{\Psi}\right\|_{F}^{2}.$
By dropping the diagonal blocks and the $(p+1)$-th row blocks of the matrix on the right-hand side, we have
$c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\lambda\sum_{i=1}^{p}\sum_{j=1,j\neq i}^{p+1}\|\bm{A}_{i}\bm{\Psi}^{(i,j)}\|_{F}^{2}\geq\frac{\lambda\sigma^{*2}_{\min}}{4}\sum_{i=1}^{p}\sum_{j=1,j\neq i}^{p+1}\|\bm{\Psi}^{(i,j)}\|_{F}^{2},$ (141)
where the second inequality follows from (58) and (71). Applying the same argument to (70) yields
$c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\frac{\lambda\sigma^{*2}_{\min}}{4}\sum_{j=1}^{p}\sum_{i=1,i\neq j}^{p+1}\|\bm{\Psi}^{(i,j)}\|_{F}^{2}.$
This, together with (141), yields
@@
$\frac{\delta}{3}\geq\frac{2\sqrt{2}c_{2}}{\sqrt{\lambda}\sigma_{\min}^{*}}\|\nabla G(\bm{W})\|_{F}\geq\left\|\bm{\Psi}-\mathrm{BlkDiag}\left(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)}\right)\right\|_{F},$ (143)
where the second inequality uses (140) and (142). This, together with $\bm{\Psi}=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}$ and Weyl’s inequality, yields for all $i\in[d_{\min}]$,
$\frac{\delta}{3}$ $\geq\left|\sigma_{i}\left(\mathrm{BlkDiag}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-y_{i}\right|$
$\geq\left|y_{i}-\sqrt{\lambda}\right|-\left|\sigma_{i}\left(\mathrm{BlkDiag}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-\sqrt{\lambda}\right|$
$\geq\delta-\left|\sigma_{i}\left(\mathrm{BlkDiag}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-\sqrt{\lambda}\right|,$
which implies $\left|\sigma_{i}(\bm{\Psi}^{(p+1,p+1)})-\sqrt{\lambda}\right|\geq{2\delta}/{3}$ for all $i\in[d_{\min}-r_{\sigma}]$. According to this, we obtain for all $i\in[d_{\min}-r_{\sigma}]$,
@@
Using (69) and (70), we have
$c_{2}\|\nabla G(\bm{W})\|_{F}\geq\left\|(\bm{\Sigma}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}})\bm{\Sigma}_{2}^{(p+1)}+\lambda\bm{\Sigma}_{2}^{(p+1)}-\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{T}_{2}^{(p+1)}\right\|_{F}.$ (146)
According to Mirsky’s inequality (see Lemma 11), we have
$\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}=\|\bm{\Sigma}_{1}^{(p+1)}-\bm{0}\|_{F}\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{1}^{*}\|_{F}\leq\|\bm{W}_{1}-\bm{W}_{1}^{*}\|_{F}\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}).$ (147)
Similarly, we have
$\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}).$ (148)
@@
$\geq\left\|\lambda\bm{T}_{2}^{(p+1)^{T}}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)^{T}}-\sqrt{\lambda}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)}\right\|_{F}-y_{1}\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}^{3},$
@@
Summing up the above two inequalities yields
@@
Repeating the argument in (149)-(4.2) with multiplying $\sqrt{\lambda}$ on (145) and $y_{1}$ on (146), we obtain
$c_{2}(y_{1}+\sqrt{\lambda})\|\nabla G(\bm{W})\|_{F}\geq$ $\left\|\lambda^{\frac{3}{2}}\bm{\Sigma}_{1}^{(p+1)^{T}}-\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\right\|_{F}$
$-\sqrt{\lambda}\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}-y_{1}\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}.$
@@
$2c_{2}(y_{1}+\sqrt{\lambda})\|\nabla G(\bm{W})\|_{F}$
$\geq$ $\ \sqrt{\lambda}\sigma_{\min}\left(\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}-\lambda\bm{I}\right)\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}+\sqrt{\lambda}\sigma_{\min}\left(\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}-\lambda\bm{I}\right)$
$\quad\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}-(\sqrt{\lambda}+y_{1})(\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}^{3}+\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}^{3})$
$\geq$ $\ \left(\frac{2\lambda\delta}{3}-(\sqrt{\lambda}+y_{1})\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}^{2}\right)\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}+\left(\frac{2\lambda\delta}{3}-(\sqrt{\lambda}+y_{1})\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}^{2}\right)\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}$
$\geq$ $\ \frac{\lambda\delta}{3}(\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}+\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}),$
where the second inequality follows from (144), and the last inequality follows from (77a), (147) and (148). This directly implies (78).
(ii) For ease of exposition, let $\bm{\Psi}:=\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\in\mathbb{R}^{d_{L}\times d_{0}}$. Let $l\in[L]$ be such that $l\in\operatorname{\arg\max}\{\sigma_{r_{\sigma}+1}(\bm{W}_{k}):{k\in[L]}\}$. Substituting (66) into (63) yields
$c_{1}\|\nabla G(\bm{W})\|_{F}$ $\geq\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}$
$\geq\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}-\sqrt{\lambda}\|\bm{\Delta}\|_{F},$
where the second inequality uses the triangular inequality and
$\bm{\Delta}:=$ $\ (\bm{H}_{l+1}-\bm{T}_{l+1})\bm{\Sigma}_{l+1}^{T}\left(\prod_{k=l+2}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)$
$\ +\bm{T}_{l+1}\bm{\Sigma}_{l+1}^{T}(\bm{H}_{l+2}-\bm{T}_{l+2})\bm{\Sigma}_{l+2}^{T}\left(\prod_{k=l+3}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)+\cdots$
$\ +\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-2}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)\bm{\Sigma}_{l-1}^{T}(\bm{H}_{l}-\bm{T}_{l}).$
Using (57) and (139), we obtain
@@
$\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}\leq\eta_{2}\|\nabla G(\bm{W})\|_{F},$
where $\eta_{2}$ is defined in (81). Using this inequality, the block structures of $\bm{\Sigma}_{l}$ and $\bm{T}_{l}$, and the fact that $l\in\operatorname{\arg\max}\{\sigma_{r_{\sigma}+1}(\bm{W}_{k}):{k\in[L]}\}$, we obtain
$\eta_{2}\|\nabla G(\bm{W})\|_{F}$ $\geq\left\|\left(\bm{\Sigma}_{l}^{(p+1)}\bm{\Sigma}_{l}^{(p+1)^{T}}\right)^{L-1}\bm{\Sigma}_{l}^{(p+1)}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(p+1)}\bm{\Sigma}_{k}^{(p+1)^{T}}\right)\right.$
$\quad\ \left.\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{(p+1)^{T}}\bm{T}_{k+1}^{(p+1)}\right)+\lambda\bm{\Sigma}_{l}^{(p+1)}\right\|$

### v3

@@ 3.2.1 Preliminary Results
For each $i\in[p]$, using Weyl’s inequality yields $|\sigma_{j}(\bm{\Sigma}_{L}^{(i)})-\sigma_{t_{i}}^{*}|\leq\|\bm{W}_{L}-\bm{W}_{L}^{*}\|\leq{\sigma_{\min}^{*}}/{2}$ for each $j\in[g_{i}]$, which implies $\|\bm{\Sigma}_{L}^{(i)}-\sigma_{t_{i}}^{*}\bm{I}\|\leq\sigma_{\min}^{*}/2$. This, together with Lemma B.4 in Appendix B.2 and $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$, yields $\|\bm{T}_{l}^{(i)}(\bm{\Sigma}_{L}^{(i)})^{2}-(\bm{\Sigma}_{L}^{(i)})^{2}\bm{T}_{l}^{(i)}\|_{F}\geq\sigma_{\min}^{*}\|\bm{T}_{l}^{(i)}\bm{\Sigma}_{L}^{(i)}-\bm{\Sigma}_{L}^{(i)}\bm{T}_{l}^{(i)}\|_{F}.$ Using this and (69), we obtain (65). ∎
3.2.2 Proof of the Error Bound
Armed with the above results, we now proceed to prove the error bound stated in 2. This is achieved in three steps. First, we show that, in a neighborhood of the critical point set, the singular values and corresponding singular vectors of the weight matrices can be controlled by $\|\nabla G(\bm{W})\|_{F}$ (see Propositions 4 and 5). Second, we establish the error bound for Problem (12) (see 3). To this end, we construct an intermediate point $\hat{\bm{W}}$ using left and right singular matrices (or suitable orthogonal matrices derived from them) of $\bm{W}$ (see (116) in the proof of 3) and show that both $\mathrm{dist}(\hat{\bm{W}},\mathcal{W}_{G})$ and $\|\bm{W}-\hat{\bm{W}}\|_{F}$ are upper bounded by $\|\nabla G(\bm{W})\|_{F}$. Finally, using Lemma 1 and 3, we prove the desired error bound of Problem (1) in 2.
We next bound the singular values and the associated singular vectors of the weight matrices in the following two propositions.
Proposition 4.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary. The following statements hold:
(i) Suppose that $L=2$ and (7) holds. Then for all $\bm{W}$ satisfying
$\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\min\left\{\frac{\delta_{\sigma}}{3},\frac{\sqrt{\lambda}}{\sqrt{3(\sqrt{\lambda}+y_{1})}}\left(\min\left\{\min_{i\in[s_{p_{Y}}]}\left|\sqrt{\lambda}-y_{i}\right|,\sqrt{\lambda}\right\}\right)^{{1}/{2}}\right\},$ (70a)
it holds for $l=1,2$ that
where $c_{3}:={6c_{2}(y_{1}+\sqrt{\lambda})}/({\lambda\min\{\min_{i\in[s_{p_{Y}}]}|\sqrt{\lambda}-y_{i}|,\sqrt{\lambda}\}}).$
(ii) Suppose that $L\geq 3$ and (8) holds. Then for all $\bm{W}$ satisfying
it holds for all $l\in[L]$ that
where
$c_{3}:=\frac{2}{\lambda}\left(c_{1}+\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3(L-1)y_{1}}{2\delta_{\sigma}\sqrt{\lambda}\sigma^{*}_{\min}}\right).$ (74)
Proof.
@@ 3.2.2 Proof of the Error Bound
Let $\bm{W}^{*}=(\bm{W}_{1}^{*},\dots,\bm{W}_{L}^{*})\in\mathcal{W}_{\bm{\sigma}^{*}}$ be such that $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})=\|\bm{W}-\bm{W}^{*}\|_{F}$. From (70a), we have $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3<\sigma_{\min}^{*}/2$ in case (i). From (72), the same bound holds in case (ii). This, together with Lemma 2 and 2, implies that (35)-(37) and (41) hold. According to $\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\delta_{\sigma}/3$, 3, and Lemma 3, there exist matrices $\bm{T}_{l}^{(i)}\in\mathcal{O}^{g_{i}}$ for all $i\in[p]$ and $\bm{T}_{l}^{(p+1)}\in\mathcal{O}^{d_{l-1}-r_{\sigma}}$ such that (48), (58), and (59) hold for all $l=2,\dots,L$, where $\bm{A}_{i}$ and $\bm{B}_{i}$ for each $i\in[p+1]$ are respectively defined in (54) and (55). Recall that (3.2.1) denotes an SVD of $\bm{W}_{l}$ and let $\bm{H}_{l}:=\bm{U}_{l-1}^{T}\bm{V}_{l}$ and $\bm{T}_{l}:=\mathrm{BlkD}(\bm{T}_{l}^{(1)},\dots,\bm{T}_{l}^{(p)},\bm{T}_{l}^{(p+1)})$ for each $l=2,\dots,L$. This, together with (48), implies
@@
(i) Suppose that $L=2$ and (7) holds. For ease of exposition, define $\bm{\Psi}:=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}\in\mathbb{R}^{d_{2}\times d_{0}}$ and we partition $\bm{\Psi}$ as
@@
where $\bm{\Psi}^{(i,j)}\in\mathbb{R}^{g_{i}\times g_{j}}$ for each $i,j\in[p]$, $\bm{\Psi}^{(p+1,j)}\in\mathbb{R}^{(d_{2}-r_{\sigma})\times g_{j}},\ \bm{\Psi}^{(i,p+1)}\in\mathbb{R}^{g_{i}\times(d_{0}-r_{\sigma})}$ for each $i,j\in[p]$, and $\bm{\Psi}^{(p+1,p+1)}\in\mathbb{R}^{(d_{2}-r_{\sigma})\times(d_{0}-r_{\sigma})}$. Using $L=2$, (54), (55), (58) and (59), we have
$c_{2}\|\nabla G(\bm{W})\|_{F}\geq\left\|(\bm{\Sigma}_{2}^{(p+1)}\bm{\Sigma}_{2}^{(p+1)^{T}})\bm{\Sigma}_{2}^{(p+1)}+\lambda\bm{\Sigma}_{2}^{(p+1)}\!\!-\!\!\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{T}_{2}^{(p+1)}\right\|_{F}.$ (78)
Using (58) again, we have
$c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\left\|(\bm{\Sigma}_{1}\bm{\Sigma}_{1}^{T})\bm{\Sigma}_{1}+\lambda\bm{\Sigma}_{1}-\sqrt{\lambda}\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{A}_{p+1}\right)\bm{\Psi}\right\|_{F}^{2}.$
By dropping the diagonal blocks and the $(p+1)$-th row blocks of the matrix on above the right-hand side, we have
[…]
Applying the same argument to (59) yields $c_{2}^{2}\|\nabla G(\bm{W})\|_{F}^{2}\geq\frac{\lambda\sigma^{*2}_{\min}}{4}\sum_{j=1}^{p}\sum_{i=1,i\neq j}^{p+1}\|\bm{\Psi}^{(i,j)}\|_{F}^{2}.$ This, together with (79), yields
@@
[…]
This, together with $\bm{\Psi}=\bm{U}_{2}^{T}\bm{Y}\bm{V}_{1}$ and Weyl’s inequality, yields for all $i\in[d_{\min}]$,
$\frac{\delta}{3}$ $\geq\left|y_{i}-\sigma_{i}\left(\mathrm{BlkD}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)\right|$
$\geq\left|y_{i}-\sqrt{\lambda}\right|-\left|\sigma_{i}\left(\mathrm{BlkD}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-\sqrt{\lambda}\right|$
$\geq\delta-\left|\sigma_{i}\left(\mathrm{BlkD}(\bm{\Psi}^{(1,1)},\dots,\bm{\Psi}^{(p+1,p+1)})\right)-\sqrt{\lambda}\right|,$
which implies $\left|\sigma_{i}(\bm{\Psi}^{(p+1,p+1)})-\sqrt{\lambda}\right|\geq{2\delta}/{3}$ for all $i\in[d_{\min}-r_{\sigma}]$. Accordingly, for all $i\in[d_{\min}-r_{\sigma}]$, we obtain
@@
$\geq\left\|\lambda\bm{T}_{2}^{(p+1)^{T}}\bm{\Sigma}_{1}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)^{T}}-\sqrt{\lambda}\bm{\Sigma}_{2}^{(p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}{\bm{\Psi}^{(p+1,p+1)}}^{T}\right\|_{F}-y_{1}\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}^{3},$
@@
Summing up the above two inequalities, together with the triangle inequality, yields
@@
Repeating the argument from (82) to (3.2.2), but this time multiplying (77) by $\sqrt{\lambda}$ and (78) by $y_{1}$, we obtain
$c_{2}(y_{1}+\sqrt{\lambda})\|\nabla G(\bm{W})\|_{F}$
$\geq$ $\left\|\lambda^{\frac{3}{2}}\bm{\Sigma}_{1}^{(p+1)^{T}}-\sqrt{\lambda}\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}\bm{\Sigma}_{1}^{(p+1)^{T}}\right\|_{F}-\sqrt{\lambda}\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}-y_{1}\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}.$
@@
$2c_{2}(y_{1}+\sqrt{\lambda})\left\|\nabla G(\bm{W})\right\|_{F}$
$\geq$ $\ \sqrt{\lambda}\left\|(\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}-\lambda\bm{I})\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}+\sqrt{\lambda}\left\|(\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}-\lambda\bm{I})\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}$
$\quad-(\sqrt{\lambda}+y_{1})\left(\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}+\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}\right)$
$\geq$ $\ \sqrt{\lambda}\sigma_{\min}\left(\bm{\Psi}^{(p+1,p+1)}\bm{\Psi}^{(p+1,p+1)^{T}}-\lambda\bm{I}\right)\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}+\sqrt{\lambda}\sigma_{\min}\left(\bm{\Psi}^{(p+1,p+1)^{T}}\bm{\Psi}^{(p+1,p+1)}-\lambda\bm{I}\right)\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}$
$\quad-(\sqrt{\lambda}+y_{1})\left(\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}^{3}+\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}^{3}\right)$
[…]
$\geq$ $\ \frac{\lambda\delta}{3}\left(\left\|\bm{\Sigma}_{1}^{(p+1)}\right\|_{F}+\left\|\bm{\Sigma}_{2}^{(p+1)}\right\|_{F}\right),$
where the last inequality follows from the definition of $\delta$, (70a), and $\|\bm{\Sigma}_{1}^{(p+1)}\|_{F}=\|\bm{\Sigma}_{1}^{(p+1)}-\bm{0}\|_{F}\leq\|\bm{\Sigma}_{1}-\bm{\Sigma}_{1}^{*}\|_{F}\leq\|\bm{W}_{1}-\bm{W}_{1}^{*}\|_{F}\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ and $\|\bm{\Sigma}_{2}^{(p+1)}\|_{F}\leq\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ due to Mirsky’s inequality (see Lemma D.8 in Appendix D). This directly implies (71).
(ii) Suppose that $L\geq 3$ and (8) holds. For ease of exposition, let $\bm{\Psi}:=\bm{U}_{L}^{T}\bm{Y}\bm{V}_{1}\in\mathbb{R}^{d_{L}\times d_{0}}$ and we still partition $\bm{\Psi}$ into the form as in (76). Let $l\in[L]$ be such that $l\in\operatorname{\arg\max}\{\sigma_{r_{\sigma}+1}(\bm{W}_{k}):{k\in[L]}\}$. Define
$\bm{\Delta}$ $\textstyle:=\left(\prod_{k=l+1}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)-\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)$
$\textstyle~=\sum_{j=l}^{L-1}\left(\prod_{k=l+1}^{j}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)(\bm{H}_{j+1}-\bm{T}_{j+1})\bm{\Sigma}_{j+1}^{T}\left(\prod_{k=j+2}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)$
$\textstyle\quad+\sum_{j=2}^{l}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{j-2}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)\bm{\Sigma}_{j-1}^{T}(\bm{H}_{j}-\bm{T}_{j})\left(\prod_{k=j}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right).$
Here, we remark that in the above summations, if the lower index exceeds the upper index, the corresponding product is interpreted as the identity matrix. Then substituting (3.2.1) into (41) yields
$c_{1}\|\nabla G(\bm{W})\|_{F}$ $\textstyle\geq\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{H}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{H}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}$
$\textstyle\geq\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}-\sqrt{\lambda}\|\bm{\Delta}\|_{F},$
where the second inequality uses the triangle inequality. Using (35) and (75), we obtain
@@
$\textstyle\left\|(\bm{\Sigma}_{l}\bm{\Sigma}_{l}^{T})^{L-1}\bm{\Sigma}_{l}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}\bm{\Sigma}_{k}^{T}\right)\bm{\Psi}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{T}\bm{T}_{k+1}\right)+\lambda\bm{\Sigma}_{l}\right\|_{F}\leq\eta_{2}\|\nabla G(\bm{W})\|_{F},$ (84)
where $\eta_{2}:=c_{1}+\left(\frac{3}{2}\sigma_{\max}^{*}\right)^{L}\frac{3(L-1)y_{1}}{2\delta_{\sigma}\sqrt{\lambda}\sigma^{*}_{\min}}$. Using this inequality, the block structures of $\bm{\Sigma}_{l}$ and $\bm{T}_{l}$, $l\in\operatorname{\arg\max}\{\sigma_{r_{\sigma}+1}(\bm{W}_{k}):{k\in[L]}\}$, and the fact that the Frobenius norm of a matrix is larger than or equal to its spectral norm, we obtain
$\eta_{2}\|\nabla G(\bm{W})\|_{F}$
$\geq$ $\left\|\left(\bm{\Sigma}_{l}^{(p+1)}\bm{\Sigma}_{l}^{(p+1)^{T}}\right)^{L-1}\bm{\Sigma}_{l}^{(p+1)}-\sqrt{\lambda}\left(\prod_{k=l+1}^{L}\bm{T}_{k}^{(p+1)}\bm{\Sigma}_{k}^{(p+1)^{T}}\right)\bm{\Psi}_{r_{\sigma}+1:d_{L},r_{\sigma}+1:d_{0}}\left(\prod_{k=1}^{l-1}\bm{\Sigma}_{k}^{(p+1)^{T}}\bm{T}_{k+1}^{(p+1)}\right)+\lambda\bm{\Sigma}_{l}^{(p+1)}\right\|$